### v1

@@ 4.2.1 Image Classification
We perform our evaluation on the ImageNet dataset (Russakovsky et al.,, 2015) and the Extended Complex Scene Saliency Dataset (ECSSD) (Shi et al.,, 2015). Thereby, we consider two established model architectures: VGG-16 (Simonyan and Zisserman,, 2015) and ResNet-50 (He et al.,, 2016), both convolutional neural networks (CNN). The ground truth masks of ECCSD are more granular than the bounding boxes of ImageNet, which contain more pixels than the actual ground truth object. Therefore, ECCSD is more suitable for measuring localization. However, for comparison and completeness, we measure sparsity, faithfulness, and robustness also on the widely used ImageNet dataset. Thereby, we use a sample of 3,900 images of its validation set covering all classes. As we must draw multiple samples to calculate the Lipschitz estimate for a single input for robustness analyses, we restrict the number of input samples to 10. The evaluation includes both of our approaches, PLRP-$\lambda$ and PLRP-$M$, with and without sparsity gain.
Our implementation of PLRP [^1] Code available at https://anonymous.4open.science/r/plrp-FD34 is based on the Zennit package (Anders et al.,, 2021). For the evaluation, we use the Quantus package (Hedström et al.,, 2023) that implements several evaluation metrics. For evaluating faithfulness, we use a feature of the iNNvestigate package (Alber et al.,, 2019) that perturbs the input $\boldsymbol{x}$ based on its explanation $e(\boldsymbol{x})$.
@@ Results
We observe that our approach creates sparser explanations, while maintaining the most relevant features. Figure 2 illustrates that the Gini Index increases drastically for all our approaches compared to the LRP baseline. Simultaneously, the RMA increases, indicating that after pruning and redistributing a higher proportion of relevance falls within the ground truth mask. In both metrics, we observe a steep increase already for small proportions ($<0.2$) of pruned relevance $p$. For both models, VGG16 and ResNet50, the simpler approach PLRP-$\lambda$ with and without sparsity gain leads to better results compared to PLRP-$M$ and the LRP baseline. We also find that the RMA starts decreasing again for higher $p$. This is not surprising as, at some point, the most decisive features might also get pruned. These findings are in line with what we observe qualitatively by visualization (see Figure 4), i.e., PLRP-$\lambda$ produces sparser explanations than PLRP-$M$ for the same pruning parameter $p$. For higher $p$, also relevance within the ground truth mask gets pruned. The entropy shows completely analogous results to the Gini Index and hence, is not displayed here.
With respect to faithfulness, except for one approach (PLRP-$M$ with and without sparsity gain and $p<0.3$), we observe slightly worse results than the baseline. Note that better results correspond to a lower AUC as we are interested in a quick drop in the prediction score. However, we observe that the higher AUC is mainly driven by less important features, i.e., we have a less steep decrease in the prediction score for later perturbed and hence, less important features. This is in line with our goal of maintaining only the most important features. In fact, for the features with the highest relevance, the prediction score drops similarly steeply as for the baseline (see Figure 3). Overall, the relative decline in faithfulness is small compared to the relative gain in sparsity and localization, especially for small $p$.
Finally, PLRP-$\lambda$ is always more robust than the LRP baseline. Only PLRP-$M$ is less robust than the baseline for few parameterizations. Therefore, the explanations of our approach can be seen as reliable and stable.
@@
Figure 2: Results for ECSSD for metrics sparsity, localization, faithfulness, and robustness for different proportions of pruned relevance mass $p$ for models VGG16 and ResNet50. For sparsity, localization, and faithfulness, the y-axis covers the whole output domain of $[0,1]$. The zoom plots focus on the actual covered output domain for better comparison of the methods.
@@
Figure 3: Results for ECSSD for faithfulness for an exemplary parameterization $p=0.15$ for models VGG16 (left) and ResNet50 (right). The prediction score $f_{c^{*}}(\boldsymbol{x})$ drops similarly steeply as for the LRP baseline for the features with the highest relevance that are perturbed first. The difference in AUC is rather driven by the less important features that are perturbed later.
Figure 4: Illustration of relevance attribution via heatmaps for LRP and PLRP with different parameterizations for the VGG16 model. Red pixels indicate positive relevance, while blue are negative.
@@
Overall, PLRP-$\lambda$ is able to increase sparsity quicker and concentrate more relevance within the ground truth mask than PLRP-$M$, i.e., there is a steeper increase for the same proportion of pruned relevance. It appears that redistributing the pruned relevance in the manner of LRP as done by PLRP-$M$ slows down the concentration of relevance. This is also in line with PLRP-$M$ sticking closer to the baseline with respect to faithfulness. Another driver for the better results of PLRP-$\lambda$ over PLRP-$M$ with respect to sparsity, localization, and robustness appears to be that PLRP-$M$ can lead to a sign flipping of a neuron’s relevance (see negative relevance for PLRP-$M$ in Figure 4(d)). PLRP-$M$ changes the proportions of how much relevance is attributed from one neuron to another (see Equation 3.5). Since this might increase the proportion of incoming negative relevance or decrease it for incoming positive relevance, respectively, the sign of the total relevance a neuron receives might switch. Furthermore, the sign of the proportions themselves might change. PLRP-$\lambda$, on the other hand, only rescales the relevance by a positive factor and hence, preserves the initial sign of the relevance score.
@@
The quantitative results for the ImageNet dataset for sparsity, faithfulness, and robustness are analogous to the results for ECSSD and can be found in the supplementary material in section A.
@@
Figure 5: Illustration of pruned explanations: Logo plots of (a) ground truth mask, (b) relevance scores for LRP, and (c) PLRP-$\lambda$ with $p=0.25$ for the model’s prediction. PLRP reduces noise and creates a sparser explanation compared to the LRP baseline. Thereby, PLRP maintains features that lie within the ground truth even if they have less relevance than the noise at other irrelevant features. Hence, it indeed differs from only applying a threshold.
@@ 5 Conclusion
In this work, we introduced PLRP, a modification of LRP that enforces sparsity directly by locally pruning the relevance propagation for the different layers. We presented two approaches to redistribute the pruned relevance mass so that the conservation property is preserved: the simpler PLRP-$\lambda$ and the more complex PLRP-$M$, which is more in line with original LRP methodology. The evaluation on the ECSSD and ImageNet datasets shows that we indeed obtain sparser explanations than the LRP baseline. The pruning leads to only a slight decrease in faithfulness, mainly driven by the least important features. By measuring localization, we show that while becoming sparser, more relevance is attributed to features within the ground truth mask. This demonstrates the efficacy of PLRP compared to LRP. In fact, both our approaches lead to noise reduction and concentrate relevance at the most important features. Qualitative evaluation by visualization supports this claim. Furthermore, our approach is similar and, depending on the parameterization, even more robust to small changes in the input than the LRP baseline. For the genomics application, we observe similar effects. We obtain sparser explanations with higher concentration of relevance on the most important features in the genome sequence, aiding in the interpretation of the model outputs. Overall, PLRP-$\lambda$ produces better results for our goal of sparsity and relevance concentration than PLRP-$M$. This is partially driven by the effect of relevance sign flipping. However, it should be investigated whether this effect can be reduced by further modification. Additionally, how to define and find an optimal parameterization for PRLP remain open questions. This might heavily depend on the underlying model but also on the “optimal” degree of sparsity for a specific task. Finally, as PLRP not only prunes the relevance attribution for the input features but also for neurons in intermediate layers, it could be used to study what concepts a model learned in deeper layers.
@@
Figure 6: Results for ImageNet dataset for metrics sparsity, faithfulness, and robustness for different proportions of pruned relevance mass $p$ for models VGG16 and ResNet50. For sparsity, and faithfulness, the y-axis covers the whole output domain of $[0,1]$. The zoom plots focus on the actual covered output domain for better comparison of the methods.
Figure 7: Results for ImageNet dataset for faithfulness for an exemplary parameterization $p=0.15$ for models VGG16 (left) and ResNet50 (right). The prediction score $f_{c^{*}}(\boldsymbol{x})$ drops similarly steeply as for the LRP baseline for the features with the highest relevance that are perturbed first. The difference in AUC is rather driven by the less important features that are perturbed later.

### v2

@@ 4.2.1 Image Classification
We perform our evaluation on the ImageNet dataset (Russakovsky et al.,, 2015) and the Extended Complex Scene Saliency Dataset (ECSSD) (Shi et al.,, 2015). Thereby, we consider two established model architectures: VGG-16 (Simonyan and Zisserman,, 2015) and ResNet-50 (He et al.,, 2016), both convolutional neural networks (CNN). The ground truth masks of ECCSD are more granular than the bounding boxes of ImageNet, which contain more pixels than the actual ground truth object. Therefore, ECCSD is more suitable for measuring localization. However, for comparison and completeness, we measure sparsity, robustness, and faithfulness also on the widely used ImageNet dataset. Thereby, we use a sample of 3,900 images of its validation set covering all classes. As we must draw multiple samples to calculate the Lipschitz estimate for a single input for robustness analyses, we restrict the number of input samples to 10. The evaluation includes both of our approaches, PLRP-$\lambda$ and PLRP-$M$, with and without sparsity gain. Thereby, we examine multiple values for the parameter $p\in[0,1)$, i.e., the proportion of pruned relevance per layer. Specifically, we consider a range from 0 to 0.95 in increments of 0.05.
Our implementation of PLRP [^1] Code available at https://gitlab.com/dacs-hpi/plrp is based on the Zennit package (Anders et al.,, 2021). For the evaluation, we use the Quantus package (Hedström et al.,, 2023) that implements several evaluation metrics. For evaluating faithfulness, we use a feature of the iNNvestigate package (Alber et al.,, 2019) that perturbs the input $\boldsymbol{x}$ based on its explanation $e(\boldsymbol{x})$.
@@ Results
We observe that our approach creates sparser explanations, while maintaining the most relevant features. Figure 2 illustrates that the Gini Index increases drastically for all our approaches compared to the LRP baseline. Simultaneously, the RMA increases, indicating that after pruning and redistributing a higher proportion of relevance falls within the ground truth mask. In both metrics, we observe a steep increase already for small proportions (<0.2<0.2) of pruned relevance $p$. For both models, VGG16 and ResNet50, the simpler approach PLRP-$\lambda$ with and without sparsity gain leads to better results compared to PLRP-$M$ and the LRP baseline. We also find that the RMA starts decreasing again for higher $p$. This is not surprising as, at some point, the most decisive features might also get pruned. These findings are in line with what we observe qualitatively by visualization (see Figure 4), i.e., PLRP-$\lambda$ produces sparser explanations than PLRP-$M$ for the same pruning parameter $p$. For higher $p$, also relevance within the ground truth mask gets pruned. The entropy shows completely analogous results to the Gini Index and hence, is not displayed here.
@@
Figure 2: Results for ECSSD for metrics sparsity, localization, and robustness for different proportions of pruned relevance mass $p$ ranging from 0 to 0.95 in increments of 0.05 for models VGG16 and ResNet50. For sparsity and localization, the y-axis covers the whole output domain of $[0,1]$. The zoom plots focus on the actual covered output domain for better comparison of the methods.
Furthermore, PLRP-$\lambda$ is always more robust than the LRP baseline, whereas PLRP-$M$ is less robust than the baseline for only a few parameterizations. Thus, we conclude that the explanations of our approach can be seen as reliable and stable.
With respect to faithfulness, except for one approach (PLRP-$M$ with and without sparsity gain and $p<0.3$), we observe slightly worse results than the baseline (see Figure 3). Note that better results correspond to a lower AUC as we are interested in a quick drop in the prediction score. However, we observe that the higher AUC is mainly driven by less important features, i.e., we have a less steep decrease in the prediction score for later perturbed and hence, less important features. This is in line with our goal of maintaining only the most important features. In fact, for the features with the highest relevance, the prediction score drops similarly steeply as for the baseline (see Figure 3 (right)). Finally, the relative decline in faithfulness is small compared to the relative gain in sparsity and localization, especially for small $p$.
@@
Figure 3: Results for ECSSD for faithfulness for different proportions of pruned relevance mass $p$ ranging from 0 to 0.95 in increments of 0.05 (left) and for an exemplary parameterization $p=0.15$ (right) for models VGG16 and ResNet50. The AUC differs only slightly compared to the LRP baseline (left). For higher $p$, the difference becomes larger as more features are pruned. For the exemplary parameterization, the prediction score $f_{c^{*}}(\boldsymbol{x})$ drops similarly steeply as for the LRP baseline for the features with the highest relevance that are perturbed first (right). The difference in AUC is driven by the less important features that are perturbed later.
@@
Overall, PLRP-$\lambda$ is able to increase sparsity quicker and concentrate more relevance within the ground truth mask than PLRP-$M$, i.e., there is a steeper increase for the same proportion of pruned relevance. It appears that redistributing the pruned relevance in the manner of LRP as done by PLRP-$M$ slows down the concentration of relevance. This is also in line with PLRP-$M$ sticking closer to the baseline with respect to faithfulness. Another driver for the better results of PLRP-$\lambda$ over PLRP-$M$ with respect to sparsity, localization, and robustness appears to be that PLRP-$M$ can lead to a sign flipping of a neuron’s relevance (see negative relevance for PLRP-$M$ in Figure 4(d)). PLRP-$M$ changes the proportions of how much relevance is attributed from one neuron to another (see Equation (3.5)). Since this might increase the proportion of incoming negative relevance or decrease it for incoming positive relevance, respectively, the sign of the total relevance a neuron receives might switch. Furthermore, the sign of the proportions themselves might change. PLRP-$\lambda$, on the other hand, only rescales the relevance by a positive factor and hence, preserves the initial sign of the relevance score.
@@
The quantitative results for the ImageNet dataset for sparsity, robustness, and faithfulness are analogous to the results for ECSSD. Hence, the experiments on ImageNet confirm the results above, thereby providing further confidence in the findings. They can be found in the supplementary material in section A.
[…]
@@
Figure 5: Illustration of pruned explanations: Logo plots of (a) ground truth mask, (b) relevance scores from LRP, and (c) PLRP-$\lambda$ with $p=0.25$ for the model’s prediction. PLRP reduces noise and creates a sparser explanation compared to the LRP baseline. Thereby, PLRP maintains features that lie within the ground truth even if they have less relevance than the noise at other irrelevant features. Hence, it indeed differs from only applying a threshold.
@@ 5 Conclusion
In this work, we introduced PLRP, a modification of LRP that enforces sparsity directly by locally pruning the relevance propagation for the different layers. We presented two approaches to redistribute the pruned relevance mass so that the conservation property is preserved: the simpler PLRP-$\lambda$ and the more complex PLRP-$M$, which is more in line with original LRP methodology. The evaluation on the ECSSD and ImageNet datasets shows that we indeed obtain sparser explanations than the LRP baseline. The pruning leads to only a slight decrease in faithfulness, mainly driven by the least important features. By measuring localization, we show that while becoming sparser, more relevance is attributed to features within the ground truth mask. This demonstrates the efficacy of PLRP compared to LRP. In fact, both our approaches lead to noise reduction and concentrate relevance at the most important features. Qualitative evaluation by visualization supports this claim. Furthermore, our approach is similar and, depending on the parameterization, even more robust to small changes in the input than the LRP baseline. For the genomics application, we observe similar effects. We obtain sparser explanations with higher concentration of relevance on the most important features in the genome sequence, aiding in the interpretation of the model outputs. Overall, PLRP-$\lambda$ produces better results for our goal of sparsity and relevance concentration than PLRP-$M$. This is partially driven by the effect of relevance sign flipping. However, it should be investigated whether this effect can be reduced by further modification. Additionally, how to define and find an optimal parameterization for PRLP remain open questions. This might heavily depend on the underlying model but also on the “optimal” degree of sparsity for a specific task. Further, as PLRP not only prunes the relevance attribution for the input features but also for neurons in intermediate layers, it could be used to study what concepts a model learned in deeper layers. Finally, as the general idea of enforcing sparsity when attributing relevance is not specific to LRP, it can also be applied to other explanation methods.
@@
Figure 6: Results for ImageNet dataset for metrics sparsity and robustness for different proportions of pruned relevance mass $p$ ranging from 0 to 0.95 in increments of 0.05 for models VGG16 and ResNet50. For sparsity, the y-axis covers the whole output domain of $[0,1]$. The zoom plots focus on the actual covered output domain for better comparison of the methods.
Figure 7: Results for ImageNet dataset for faithfulness for different proportions of pruned relevance mass $p$ ranging from 0 to 0.95 in increments of 0.05 (left) and for an exemplary parameterization $p=0.15$ (right) for models VGG16 and ResNet50. The AUC differs only slightly compared to the LRP baseline (left). For higher $p$, the difference becomes larger as more features are pruned. For the exemplary parameterization, the prediction score $f_{c^{*}}(\boldsymbol{x})$ drops similarly steeply as for the LRP baseline for the features with the highest relevance that are perturbed first (right). The difference in AUC is driven by the less important features that are perturbed later..